OpenStudy (anonymous):

You are given a pair of equations, one representing a supply curve and the other representing a demand curve, where p is the unit price for x items. and 50p+x-1148=0, 86p-x-28=0, Identify which is the supply curve and demand curve and the appropriate domain. Put the domains in interval notation . For type a.infinity . For more than one interval use a U to represent a "union". Domain of the supply curve Domain of the demand curve b. Determine the market equilibrium. Equilibrium: x = p = c. Determine the revenue function. Revenue function R(x)= d. Determine the revenue at market equi

OpenStudy (bahrom7893):

okay lol i just figured out how to do this.. hi again lol

OpenStudy (bahrom7893):

Demand curve always has a negative slope, while supply curve always has a positive slope.

OpenStudy (bahrom7893):

Now solve for p in both equations to see the slope: 1)50p+x-1148=0, subtract x from both sides: 50p-1148 = -x, add 1148 to both sides: 50p = -x + 1148, divide both sides by 50: p = (-1/50)x + (1148/50)

OpenStudy (bahrom7893):

now that's in the form of y = Mx+b, where M is the slope. In this case M = (-1/50), which is negative. So this is the demand curve (well line in this case).

OpenStudy (bahrom7893):

repeat the same thing for the other equation, it should be supply

OpenStudy (anonymous):

okay

OpenStudy (bahrom7893):

let me work this out: 86p-x-28=0 86p = x+28 p = (1/86)x + (14/43) - supply curve

OpenStudy (bahrom7893):

Okay so we have: Supply: p = (1/86)x + (14/43) Demand: p = (-1/50)x + 22.96

OpenStudy (bahrom7893):

Domain of both of these are: (-infinity; infinity) since x can be any real number, these are just straight continuous lines. By the way, I'd appreciate it if you don't mention that I go to the same school that u go to. U can tell them where u go, but i already feel like most ppl on here know too much about me haha.

OpenStudy (bahrom7893):

Looking up what market equilibrium is...

OpenStudy (bahrom7893):

Ohh market equilibrium is when supply = demand, so set the two equations equal to each other

OpenStudy (bahrom7893):

(1/86)x + (14/43) = (-1/50)x + 22.96 solve for x

OpenStudy (bahrom7893):

only professors from our school are going to give you retarded numbers on webworks.. freakin hate that...

OpenStudy (bahrom7893):

(1/86)x + (14/43) = (-1/50)x + 22.96 x((1/86)+(1/50)) = 22.96 - (14/43) (34/1075)x = 24332/1075 hmm the numbers do workout nicely haha

OpenStudy (anonymous):

professor is very fast with her lecture...

OpenStudy (bahrom7893):

multiply everything out by 1075: 34x = 24332 divide everything by 34: x = 24332/34 = 12166/17

OpenStudy (bahrom7893):

Nevermind, the numbers are ugly: x = 12166/17 p = (1/86)(12166/17)+ (14/43) = 147/17

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

Revenue = x*demand = x*[(-1/50)x + 22.96] = (-1/50)x^2 + 22.96x

OpenStudy (bahrom7893):

u might have to put the asterix, sometimes webworks is too dumb to get that it's multiplication: Revenue function: (-1/50)*x^2 + 22.96*x

OpenStudy (bahrom7893):

For revenue at market equilibrium just plug in x = 12166/17 into revenue function.

OpenStudy (bahrom7893):

R(12166/17) = (-1/50)*(12166/17)^2 + 22.96*(12166/17) if u just copy paste this, webworks should accept it: (-1/50)*(12166/17)^2 + 22.96*(12166/17)

OpenStudy (anonymous):